# Fourier neural operator for large eddy simulation of compressible Rayleigh-Taylor turbulence

### v1

@@ V A POSTERIORI TEST
In this section, the a posteriori tests for ILES and LES using the FNO, VGM, and DSM models of 3D compressible RT turbulence are performed. Fig. 4 shows the transient contours of the concentration of heavy fluid at $t/\tau=2.7$ in ground truth flow field. We observe that the overall structures in different cases are similar to each other. Our focus is on predicting the statistics of the flow field and the spatial structures in the a posteriori tests. All statistical results are ensemble averages of the 5 cases, which are independently generated with respective to the data used in the training of the FNO model.
@@ V.1 Model convergence and generalization on longer time
Fig. 7 displays the predictions of the FNO models at epochs 36, 38, 40, 42, and 44, compared to the filtered ground truth (fGT): Fig. 7 (a) shows the evolution of bubble and spike heights, where $H_{B,0.1}$ or $H_{S,0.1}$ represents the height of bubble or spike calculated based on a $10\%$ threshold of concentration field, where the mean concentration is 0.9 or 0.1;Cook and Dimotakis (2001); Livescu et al. (2009); Luo and Wang (2022) Fig. 7 (b) provides the concentration profiles $\langle c\rangle_{xy}$ at time $t/\tau=2.1$. We observe that the FNO models can accurately predict the evolution of mixing heights over time and the change in average concentration with height very well as compared to the filtered ground truth (fGT) in the a posteriori test. The prediction results of five FNO models overlap with each other, indicating the convergence of the FNO models. The statistics demonstrate that the FNO model trained with the data at $t^{*}\geq 2.0$ exhibits time generalization in predicting the concentration field and can be used for long-term predictions. In the compressible RT turbulence at current parameters, since the profile of the average concentration is nearly self-similar Luo and Wang (2021), the mixing heights predicted by the FNO models at $t^{*}\geq 2.0$ still match well with the filtered ground truth. Fig. 5 (b) presents the contours of the concentration of heavy fluid at different times $t^{*}$ predicted by the FNO model at epoch 40. It can be observed that the contours of the concentration predicted by the FNO model are essentially the same as those from filtered ground truth. But errors gradually accumulate over time. At the later stages, differences between the contours predicted by the FNO model and those from filtered ground truth emerge, although the basic flow structures remain similar. There are no non-physical phenomena observed in the flow field at $t^{*}\geq 2.0$.
@@
Fig. 8 illustrates the time evolution of spatially averaged kinetic energy $K_{z}(0)=\langle\frac{\bar{\rho}}{2}\tilde{u}_{i}\tilde{u}_{i}\rangle_{xy}$ at $z=0$ plane and the kinetic energy $KE=\langle\frac{\bar{\rho}}{2}\tilde{u}_{i}\tilde{u}_{i}\rangle$ in the computational domain predicted by FNO models at epochs 36, 38, 40, 42, and 44 in the a posteriori test. The kinetic energy predicted by the five models nearly overlaps with each other, confirming the convergence of the FNO models. The time evolution of the kinetic energy $K_{z}(0)$ at $z=0$ can be roughly divided into three stages: a rapid growth stage at $t^{*}\leq 1.0$, a gradual decrease stage at $1.0<t^{*}\leq 2.0$, and a steady stage at $t^{*}>2.0$. Surprisingly, the FNO models perform well in predicting at all three stages, even though the filtered ground truth (fGT) data at the third stage was not used during training. This suggests that the FNO model has learned certain physical principles at $z=0$. However, for averaged kinetic energy $KE$ in the computational domain, the FNO models predict results close to filtered ground truth up to $t^{*}\leq 2.2$, but deviate from filtered ground truth thereafter. This deviation may mainly be due to the asymmetric exponential initial density stratification, resulting in poorer self-similarity of the density field. Although there is no significant rapid growth in the absolute errors of the density and velocity fields at $t^{*}\geq 2.0$, statistical quantities derived from these two fundamental quantities exhibit deviations with filtered ground truth. The region outside the mixing layer is in a hydrostatic state, which makes it difficult to learn the evolution of the flow field at $t^{*}>2.0$ using the FNO model trained with data $t^{*}\leq 2.0$.
@@ V.2 Comparison between FNO model and traditional LES methods
Fig. 14 displays the time evolution of the spatially averaged velocity divergence $\left\langle\theta\right\rangle_{m}$ and its rms value $\theta^{\prime}$ in the mixing region. Overall, the FNO model significantly outperforms the ILES, VGM, and DSM models in predicting velocity divergence. The mean velocity divergence $\left\langle\theta\right\rangle_{m}$ fluctuates towards a positive value, consistent with previous studiesLuo and Wang (2021, 2022). The FNO model fully captures this trend of fluctuation, while the mean velocity divergence predicted by traditional LESs shows little fluctuation with time. Similar to the $\omega^{\prime}(0)$, the rms value of velocity divergence $\theta^{\prime}$ in the mixing layer also peaks around $1.3t/\tau$. The prediction of $\theta^{\prime}$ by the FNO model coincides with filtered ground truth in the early stages but exhibits some deviation in the later stages. In contrast to the prediction of $\omega^{\prime}(0)$, the FNO model overestimates $\theta^{\prime}$. On the other hand, predictions of $\theta^{\prime}$ by ILES, VGM, and DSM models show significant discrepancies from filtered ground truth, with their rms values nearly ten times lower than filtered ground truth in the later stages. The flow field simulated by the traditional LES method tends to be in a state with zero velocity divergence, which is inconsistent with the ground truth. It is worth noting that in the initial stages of LESs, statistical quantities of velocity divergence have large values, which is caused by the subsequent calculation using coarse grids after reading the filtered ground truth flow field at $t/\tau=0.24$. The enlarged image of $\left\langle\theta\right\rangle_{m}$ at $0.25\leq t/\tau\leq 0.35$ is shown in the inset. The diffusive mixing of the two unequal-density fluids leads to large velocity divergence. Luo and Wang (2021, 2022); Cook and Dimotakis (2001) Due to insufficient grid resolution, there is a significant discretization error for velocity divergence at the interface. It is worth noting that even though the FNO model was trained using data at $t/\tau\leq 2.0$, satisfactory results were obtained even beyond $t/\tau\geq 2.0$ in the a posteriori test, including statistical quantities of velocity and vorticity at $z=0$, as well as velocity divergence in the mixing layer.
@@ V.3 Generalization on higher Reynolds numbers
[…]
@@ VI CONCLUSIONS
Subsequent work will be focused on improving the generalization ability of the FNO model for different initial conditions, stratification parameters, and Atwood numbers. In addition, neural network methods have experienced rapid development in recent years, and more advanced models can be attempted for predicting compressible RT turbulence.

### v2

@@ V A POSTERIORI TEST
In this section, the a posteriori tests for ILES and LES using the FNO, VGM, and DSM models of 3D compressible RT turbulence are performed. Fig. 4 shows the transient contours of the concentration of heavy fluid at $t/\tau=2.7$ in ground truth flow field. We observe that the overall structures in different cases are similar to each other. Our focus is on predicting the statistics of the flow field and the spatial structures in the a posteriori tests. All statistical results are ensemble averages of the 5 cases, which are independently generated with respective to the data used in the training of the FNO model. The first flow field obtained by the FNO model is at $t/\tau=0.3$. The flow field at $t/\tau\leq 0.24$ is the ground truth flow field.
@@ V.1 Model convergence and generalization on longer time
Fig. 7 displays the predictions of the FNO models at epochs 36, 38, 40, 42, and 44, compared to the filtered ground truth (fGT): Fig. 7 (a) shows the evolution of bubble and spike heights, where $H_{B,0.1}$ or $H_{S,0.1}$ represents the height of bubble or spike calculated based on a $10\%$ threshold of concentration field, where the mean concentration is 0.9 or 0.1;Cook and Dimotakis (2001); Livescu et al. (2009); Luo and Wang (2022) Fig. 7 (b) provides the concentration profiles $\langle c\rangle_{xy}$ at time $t/\tau=2.1$. We observe that the FNO models can accurately predict the evolution of mixing heights over time and the change in average concentration with height very well as compared to the filtered ground truth (fGT) in the a posteriori test. The prediction results of five FNO models overlap with each other, indicating the convergence of the FNO models. The statistics demonstrate that the FNO model trained with the data at $t^{*}\geq 2.0$ exhibits time generalization in predicting the concentration field and can be used for long-term predictions.
In the compressible RT turbulence at current parameters, since the profile of the average concentration is nearly self-similar Luo and Wang (2021), the mixing heights predicted by the FNO models at $t^{*}\geq 2.0$ still match well with the filtered ground truth. Fig. 5 (b) presents the contours of the concentration of heavy fluid at different times $t^{*}$ predicted by the FNO model at epoch 40. It can be observed that the contours of the concentration predicted by the FNO model are essentially the same as those from filtered ground truth in the early stages. But errors gradually accumulate over time. At the later stages, differences between the contours predicted by the FNO model and those from filtered ground truth emerge, although the basic flow structures remain similar. At $t^{*}\geq 2.0$, the flow field predicted by the FNO model is close to the result of solving NS equations. Qualitatively, the flow field does not exhibit any phenomena that are inconsistent with the ground truth results of RT turbulence.
@@
Fig. 8 illustrates the time evolution of spatially averaged kinetic energy $K_{z}(0)=\langle\frac{\bar{\rho}}{2}\tilde{u}_{i}\tilde{u}_{i}\rangle_{xy}$ at $z=0$ plane and the kinetic energy $KE=\langle\frac{\bar{\rho}}{2}\tilde{u}_{i}\tilde{u}_{i}\rangle$ in the computational domain predicted by FNO models at epochs 36, 38, 40, 42, and 44 in the a posteriori test. The kinetic energy predicted by the five models nearly overlaps with each other, confirming the convergence of the FNO models. The time evolution of the kinetic energy $K_{z}(0)$ at $z=0$ can be roughly divided into three stages: a rapid growth stage at $t^{*}\leq 1.0$, a gradual decrease stage at $1.0<t^{*}\leq 2.0$, and a steady stage at $t^{*}>2.0$. Surprisingly, the FNO models perform well in predicting at all three stages, even though the filtered ground truth (fGT) data at the third stage was not used during training. This suggests that the FNO model has learned certain physical principles at $z=0$. However, for averaged kinetic energy $KE$ in the computational domain, the FNO models predict results close to filtered ground truth up to $t^{*}\leq 2.2$, but deviate from filtered ground truth thereafter. This deviation may mainly be due to the asymmetric exponential initial density stratification, resulting in poorer self-similarity of the density field. The region outside the mixing layer is in a hydrostatic state, which makes it difficult to learn the evolution of the flow field at $t^{*}>2.0$ using the FNO model trained with data $t^{*}\leq 2.0$.
Although FNO is purely data-driven, it exhibits good generalization performance on longer-time predictions in the decaying HIT.Li et al. (2023b) Here, we test generalization performance of FNO in RT turbulence and observe that the predictions are satisfactory. The generalization ability of the FNO model benefits from the self similarity in the later stage of RT turbulence. However, for compressible RT turbulence, self similarity exists within a certain parameter and time range.Luo and Wang (2021, 2022); Gauthier (2017) The generalization performance is good for the current parameters and time range. For the current statistical results, only kinetic energy (Fig. 8 (b)) has deviations at $t>2$, but this degree of deviation is acceptable.
@@ V.2 Comparison between FNO model and traditional LES methods
Fig. 14 displays the time evolution of the spatially averaged velocity divergence $\left\langle\theta\right\rangle_{m}$ and its rms value $\theta^{\prime}$ in the mixing region. Overall, the FNO model significantly outperforms the ILES, VGM, and DSM models in predicting velocity divergence. The mean velocity divergence $\left\langle\theta\right\rangle_{m}$ fluctuates towards a positive value, consistent with previous studiesLuo and Wang (2021, 2022). The FNO model fully captures this trend of fluctuation, while the mean velocity divergence predicted by traditional LESs shows little fluctuation with time. Similar to the $\omega^{\prime}(0)$, the rms value of velocity divergence $\theta^{\prime}$ in the mixing layer also peaks around $1.3t/\tau$. The prediction of $\theta^{\prime}$ by the FNO model coincides with filtered ground truth in the early stages but exhibits some deviation in the later stages. In contrast to the prediction of $\omega^{\prime}(0)$, the FNO model overestimates $\theta^{\prime}$. On the other hand, predictions of $\theta^{\prime}$ by ILES, VGM, and DSM models show significant discrepancies from filtered ground truth, with their rms values nearly ten times lower than filtered ground truth in the later stages. The flow field simulated by the traditional LES method tends to be in a state with zero velocity divergence, which is inconsistent with the ground truth. It is worth noting that in the initial stages of LESs, statistical quantities of velocity divergence have large values, which is caused by the subsequent calculation using coarse grids after reading the filtered ground truth flow field at $t/\tau=0.24$. The enlarged image of $\left\langle\theta\right\rangle_{m}$ at $0.25\leq t/\tau\leq 0.35$ is shown in the inset. The diffusive mixing of the two unequal-density fluids leads to large velocity divergence. Luo and Wang (2021, 2022); Cook and Dimotakis (2001) Due to insufficient grid resolution, there is a significant discretization error for velocity divergence at the interface.
It is worth noting that even though the FNO model was trained using data at $t/\tau\leq 2.0$, satisfactory results were obtained even beyond $t/\tau\geq 2.0$ in the a posteriori test, including statistical quantities of velocity and vorticity at $z=0$, as well as velocity divergence in the mixing layer.
@@ V.3 Generalization on higher Reynolds numbers
$Re_{\lambda,z}(z)=Re\frac{\langle\rho\rangle_{xy}\lambda_{3}\left[\langle u_{3}^{\prime 2}\rangle_{xy}\right]^{1/2}}{\mu},\quad Re_{\lambda,xy}(z)=Re\frac{\langle\rho\rangle_{xy}\lambda_{i}\left[\langle u_{i}^{\prime 2}\rangle_{xy}\right]^{1/2}}{\mu}\ (\rm sum\ on\ \textit{i}=1,2),$ (27)
where $u_{i}^{\prime}=u_{i}-\langle{u}_{i}\rangle_{xy}$ is the fluctuation velocity. The three Taylor microscales are given by $\lambda_{i}(z)=\left[\frac{\langle u_{i}^{\prime 2}\rangle_{xy}}{\langle\left(\partial u_{i}^{\prime}/\partial x_{i}\right)^{2}\rangle_{xy}}\right]^{1/2}\ (\rm no\ sum\ on\ \textit{i},\>and\>\textit{i}=1,2,3)$.
Fig. 18 depicts the time evolution of horizontal and vertical Taylor Reynolds numbers at $z=0$ for Reynolds numbers $Re=10000$ and $30000$. The Taylor Reynolds number at $Re=30000$ is more than twice that at $Re=10000$. Similar to the rms values of velocities, $Re_{\lambda,z}(0)$ reaches its maximum value earlier than $Re_{\lambda,xy}(0)$. At $Re=10000$, the FNO model accurately predicts the evolution of both horizontal and vertical Taylor Reynolds numbers, outperforming traditional LESs, especially concerning the $Re_{\lambda,z}(0)$. For the case of high Reynolds number $Re=30000$, the FNO model continues to provide reasonably accurate predictions of the Taylor Reynolds numbers. It exhibits the best performance in predicting the trends of Taylor Reynolds numbers over time. The predictions of Taylor Reynolds numbers by the DSM model at both Reynolds numbers are unsatisfactory.
@@ VI CONCLUSIONS
The generalization performance and reliance on DNS data for training are obvious drawbacks of FNO models, which are also present in many machine learning methods. How to improve the generalization performance of the FNO method and reduce its reliance on DNS data is also what we need to do in the future. In addition, neural network methods have experienced rapid development in recent years, and more advanced models can be attempted for predicting compressible RT turbulence.
@@ Acknowledgements.
Appendix A The comparison between $512^{2}\times 1024$, $256^{2}\times 512$, and $128^{2}\times 256$ grids
We compare some statistical results in numerical simulations with $512^{2}\times 1024$, $256^{2}\times 512$, and $128^{2}\times 256$ grids in Fig. 22, and differences are not significant. $128^{2}\times 256$ grid resolution can reliably predict the compressible RT turbulence and is enough to properly capture the low-order statistics considered in this work.
Figure 22: Statistics in the numerical simulations with $512^{2}\times 1024$, $256^{2}\times 512$, and $128^{2}\times 256$ grids: (a) the evolutions of spike and bubble heights $H_{S/B,0.05}$, (b) the time evolution of spatially averaged kinetic energy in the computational domain $KE$, (c) the time evolution of the rms value of temperature at $z=0$, $T^{\prime}(0)$.